# Nucleus and Postperihelion Activity of Interstellar Object 3I/ATLAS Observed by Hubble Space Telescope

### v1

@@ Abstract
We report the successful detection of the nucleus of interstellar object 3I/ATLAS, achieved by applying the nucleus extraction technique to our Hubble Space Telescope (HST) observations from December 2025 to January 2026. The product of the V-band geometric albedo, $p_{V}$, with the physical cross-section of the nucleus is $0.22\pm 0.07$ km2, which corresponds to an effective radius of $1.3\pm 0.2$ km if assuming $p_{V}=0.04$, as is typical for cometary nuclei in the solar system. This size is in agreement with our estimate derived from the reported nongravitational effect and activity of the interstellar object. If the measured photometric variations are solely due to the rotation of an aspherical nucleus, the axis ratio must be $2:1$ or greater, and the rotation period $\gtrsim\!1$ hr. Leveraging the range of covered phase angles, we identified a significant opposition surge of $\sim\!0.2$ mag with a width of $3\arcdeg\pm 1\arcdeg$, which may include concurrent contributions from orbital plane crossing and tail projection, and determined a linear phase slope of $0.026\pm 0.006$ mag degree-1 for the coma dust. Compared to the preperihelion brightening trend, 3I faded more rapidly on the outbound leg, following an activity index of $4.5\pm 0.3$, not unusual in the context of solar system comets. This activity asymmetry is further corroborated by a postperihelion coma surface brightness profile that is significantly shallower than its preperihelion counterpart. From the statistics, we infer that multiple interstellar objects resembling 3I likely went undetected even before the discovery of 1I/‘Oumuamua.
@@ I Introduction
3I/ATLAS, formerly designated C/2025 N1 (ATLAS), was discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) sky survey on 1 July 2025. Serendipitous prediscovery observations were quickly identified (e.g. A. D. Feinstein et al., 2025; J. Martinez-Palomera et al., 2025; Q. Ye et al., 2025), along with numerous followup observations, leading to the definitive identification of the object as the third interstellar object to intrude into our solar system (L. Denneau et al., 2025). Its trajectory is significantly more hyperbolic with respect to the Sun than those of the previous two interstellar objects, 1I/‘Oumuamua and 2I/Borisov, with an eccentricity of $e=6.14$. The closest approach of the object to the Sun occurred on 29 October 2025 at a perihelion distance of $q=1.4$ au, implying an excess speed of $\sim\!58$ km s-1, which likely suggests a kinematic age of 3-11 Gyr for 3I, older than the previous interstellar objects (A. G. Taylor & D. Z. Seligman, 2025). The object is statistically unlikely to have approached any star as closely as the current approach to the Sun (D. Jewitt & J. Luu, 2025).
Numerous observations of 3I have been conducted using both ground-based facilities and spacecraft, spanning a wide range of wavelengths and utilising various techniques including imaging, spectroscopy, and polarimetry. In brief the interstellar object was reported to exhibit a steady growth of activity as it approached the Sun (e.g. M. R. Combi et al., 2025; D. Jewitt & J. Luu, 2025; T. Santana-Ros et al., 2025), with peculiar dust scattering properties (Z. Gray et al., 2025; S. Choi et al., 2026), slightly enriched volatile content, and distinct outgassing behaviour (e.g. M. A. Cordiner et al., 2025; D. Hutsemékers et al., 2025; N. X. Roth et al., 2025). Its activity was dominated by CO2 at heliocentric distances $r_{\rm H}\gtrsim 3$ au, which was later overtaken by H2O sublimation when closer to the Sun (e.g. M. R. Combi et al., 2025; M. A. Cordiner et al., 2025; Z. Xing et al., 2025; B. Yang et al., 2025). Conversely, there has been no successful observation of the nucleus of 3I to date, primarily because of hindrance by its activity, which began prior to the earliest available observations (e.g. Q. Ye et al., 2025). Our previous attempt to extract the nucleus signal of 3I from preperihelion Hubble Space Telescope (HST) observations in July 2025 yielded an upper limit of $\sim\!2.8$ km for the nucleus radius (D. Jewitt et al., 2025). Nevertheless, characterising the radius of the nucleus is scientifically pivotal, as an essential input for physical models and a constraint on the population of interstellar objects as a whole. In this letter, we report the first successful detection of the nucleus of 3I using postperihelion optical imaging from our HST program.
@@ III.1 Conservative Constraint
The bright and centrally condensed coma of 3I precluded direct measurement of the nucleus. We therefore attempted to isolate and remove the coma contribution to the photometry of the nucleus in different ways. The simplest method is to treat the coma contiguous to the photocentric region as the foreground (background in photometry), which inevitably underestimates the contribution of the coma in the near-nucleus region, thereby yielding an upper limit for the nucleus signal.
@@
The apparent magnitudes are functions of heliocentric distance ($r_{\rm H}$), HST-centric distances ($\Delta$), and phase angle ($\alpha$), all of which evolved over the course of our HST observations. We normalised both distances to $r_{\oplus}=1$ au and corrected for phase effects by assuming a linear phase model with a slope of $\beta_{\alpha}=0.04\pm 0.02$ mag deg-1 (e.g. P. L. Lamy et al., 2004; M. M. Knight et al., 2024) to compute the absolute magnitude:
@@
where $m_{{\rm n},V}$ is the apparent magnitude of the nucleus, $r_{\rm H}$ and $\Delta$ are expressed in au, and $\alpha$ is phase angle in degrees. We thereby determined a lower bound of $H_{{\rm n},V}\gtrsim 15$ across the five HST epochs for the nucleus of the interstellar object, consistent with $H_{{\rm n},V}\gtrsim 15.4$ from the preperihelion HST measurement (D. Jewitt et al., 2025).
@@ III.2 Nucleus Extraction
$\Sigma\left(\rho,\theta\right)=k_{\rm n}\mathcal{P}\left(\rho,\theta\right)+\left[k_{\rm c}\left(\theta\right)\left(\dfrac{\rho}{\rho_{0}}\right)^{-\gamma\left(\theta\right)}\right]\ast\mathcal{P}\left(\rho,\theta\right),$ (3)
where the total surface brightness is a function of both the projected distance from the photocenter, $\rho$, and the azimuthal angle, $\theta$. The first term on the right-hand side of Equation (3) represents the nucleus signal, modelled as the volume-normalised point-spread function (PSF, denoted $\mathcal{P}$) of the WFC3 camera using the TinyTim package (J. E. Krist et al., 2011) and scaled by a constant factor $k_{\rm n}$. The second term represents the contribution from the coma, approximated by an azimuth-dependent power-law model. The nucleus + coma model is convolved with the instrument PSF, with the symbol $\ast$ denoting the convolution operator. The pre-convolution surface brightness profile of the coma is parametrised by scaling factor $k_{\rm c}$ and logarithmic surface brightness gradient $\gamma$, both as functions of the azimuthal angle. A spherically symmetric coma in steady state has $\gamma=1$, whereas solar radiation pressure steepens the gradient to $\gamma=1.5$ in the sunward direction (D. C. Jewitt & K. J. Meech, 1987). We adopted a normalisation distance of $\rho_{0}=1$ pixel to ensure that both scaling factors share the same units.
@@
We next performed PSF photometry on the residual signal with psffit from Buie’s IDL library. Results for each HST visit are listed in Table 2, while comparisons of the observed and modelled radial brightness profiles and the apparent magnitude time series are shown in Figures 3 and 4, respectively. We were fully aware that the extracted nucleus flux possibly fell inside the caution zone where results can be biased by the nucleus extraction technique (M.-T. Hui & J.-Y. Li, 2018). To investigate the reliability of the results, we varied the azimuthal smoothing window width, image subsampling factor, and the fitted annular region within the ranges suggested by M.-T. Hui & J.-Y. Li (2018), and we even switched to aperture photometry on the residual feature using an aperture radius corresponding to its FWHM so as to better exclude jet-like features in the near-nucleus region. We found the results to be consistent within the noise level regardless of the parameter choices. Furthermore, applying the method described in Section III.1 to the whole image sequence yielded the same temporal trends as those shown in Figure 4, albeit with a systematic brightness offset due to contamination from the coma. Therefore, we conclude that the nucleus extraction technique successfully revealed the nucleus of 3I. Our previous attempt using preperihelion HST observations (D. Jewitt et al., 2025) was likely unsuccessful both as a result of lower spatial resolution at a greater (3.0 au) geocentric distance and a steeper preperihelion radial brightness profile of the coma.
Figure 4 shows tentative evidence for brightness variations, consistent with the rotation of an aspherical nucleus and/or with activity fluctuations in the near-nucleus coma. Unfortunately, our data are too sparsely sampled to permit any useful estimate of the nucleus rotation period, except that it is likely $\gtrsim\!1$ hour. If rotation is the pure cause of the brightness variations we use the observed magnitude range, $\gtrsim\!0.8$ mag, to estimate the nucleus axis ratio projected into the plane of the sky of $\gtrsim\!2:1$. However, this axis ratio could be lower, and even potentially all the way down to unity, if a significant portion of the variability stems from dust activity. Averaging these measurements yields a nucleus radius $R_{\rm n}=1.3\pm 0.2$ km for 3I, much larger than the nuclei of 1I/‘Oumuamua (highly aspherical, but effective radius $\sim\!0.08$ km) and 2I/Borisov ($\sim\!0.4$ km) (D. Jewitt & D. Z. Seligman, 2023, and citations therein).
@@
Figure 2: Detection of the nucleus of 3I (third panel from the left) via subtraction of the best-fit coma model (second panel) from the observed HST images (first panel) for each HST visit (indicated at right). For illustrative purposes, the shown images are median-combined from individual exposures within each visit and displayed with a logarithmic intensity stretch. In each row, the red and magenta arrows indicate local north and east, respectively, in the J2000 equatorial coordinate system, with the projected antisolar direction and the negative heliocentric velocity of 3I represented by the yellow and cyan arrows, respectively. The horizontal white bar near the bottom marks a scale of 1″ in apparent length.
@@
Figure 4: Temporal apparent magnitude variations of the nucleus of 3I across the HST visits. The horizontal dashed line in each panel marks the mean value computed from the respective visit.
Table 3: Comparisons of Nucleus Size Estimates of 3I/ATLAS footnotetext: The collimation coefficient of activity was assumed to be $\kappa=0.5$. For alternative values of $\kappa\in\left(0,1\right)$, the size estimate scales as $\propto\kappa^{1/3}$ per Equation (4).footnotetext: To ensure consistency, estimates originally derived by other authors assuming $p_{V}=0.05$ have been rescaled to match our adopted value of $p_{V}=0.04$.
@@ IV.1 Nucleus Size Inferred from Nongravitational Effect
The nongravitational effect of 3I allowed us to estimate its nucleus size from an independent approach, which in turn can be compared with the results derived from our HST observations. The orbital solution of 3I by JPL Horizons demonstrates a statistically significant detection of a nongravitational effect, presumably due to anisotropic mass loss from the object. Using an inverse square law for the nongravitational force model, i.e. $g_{\rm NG}\propto r_{\rm H}^{-2}$, the solution yields the radial, transverse, and normal (RTN) components of the nongravitational parameters, originally defined by B. G. Marsden et al. (1973), to be $A_{1}=\left(+4.5\pm 0.1\right)\times 10^{-8}$ au day-2, $A_{2}=\left(+1.7\pm 0.1\right)\times 10^{-8}$ au day-2, and $A_{3}=\left(-6.0\pm 0.3\right)\times 10^{-9}$ au day-2, respectively. [^2] Retrieved on 26 January 2026. Based on the law of conservation of momentum and assuming a spherical nucleus, the nucleus radius can be inferred from the nongravitational effect and the outgassing mass loss using the formulation by M.-T. Hui et al. (2020):
@@
We evaluated the nongravitational effect of 3I around the time of perihelion when its outgassing activity and the resulting nongravitational acceleration were most pronounced. At $r_{\rm H}=1.401$ au, the interstellar object exhibited significant water production at a rate of $Q_{\rm gas}=\left(3.2\pm 1.0\right)\times 10^{29}$ s-1, which subsequently declined during the outbound leg (M. R. Combi et al., 2025).  [^3] At the time of writing, H. Tan et al. (2026) reported water production rates up to an order of magnitude lower than M. R. Combi et al. (2025), despite using the same dataset. While this difference deserves to be investigated, applying the correlation from L. Jorda et al. (2008) to the near-perihelion magnitudes of 3I reported by Q. Zhang & K. Battams (2025) yielded water production rates of 3I in good agreement with M. R. Combi et al. (2025) at similar heliocentric distances, albeit preperihelion, and so we do not adopt the results of H. Tan et al. (2026). Setting $\mathscr{U}_{\rm gas}=18$ for water, $\kappa=0.5$ to approximate hemispherical emission, and adopting the empirical law for outflow speed of gas from A. L. Cochran & D. G. Schleicher (1993) yielding $v_{\rm gas}\approx 0.7$ km s-1, we obtained the nucleus radius of 3I from Equation (4) to be $1.5\pm 0.1$ km, in excellent agreement with our results in Section III.2 via the nucleus extraction technique, thereby solidifying our nucleus size estimate.
@@
Here the formal uncertainty was properly propagated from the covariance matrix of the RTN nongravitational parameters provided by Horizons and the reported error in the water production rate in M. R. Combi et al. (2025). However, we emphasise that this uncertainty is likely a significant underestimate, as it does not account for the systematic uncertainty in the unknown activity collimation coefficient. The maximum nucleus size permitted by the detected nongravitational acceleration occurs under a fully collimated mass-loss scenario, i.e. $\kappa=1$, which yields $R_{\rm n}\leq 1.9\pm 0.2$ km. Furthermore, this size estimate should also be considered as an upper limit to the radius if it should be the case that a significant part of the measured water production is due to the sublimation of icy grains in the coma.
@@
Our inferred nucleus size estimate is notably much larger than earlier results by T. M. Eubanks et al. (2025) and J. C. Forbes & H. Butler (2025), despite using the same methodology. We posit that earlier solutions of the nongravitational parameters were less reliable due to the shorter observed arc, where potential biases, such as tailward offsets in astrometric measurements, could be more easily absorbed into the orbital fits. Furthermore, the applied nongravitational force model might have been overly simplified and therefore failed to closely approximate earlier epochs when outgassing of 3I was still weak and the nongravitational effect was far less insignificant.
In Table 3, we compare our nucleus size estimates of 3I with earlier results. Although the nucleus extraction technique yielded results falling within the caution zone, we remain confident that our nucleus size estimate through this technique is trustworthy because of the consistency between these two independent methods.

### v2

@@ Abstract
We report the detection of the nucleus of interstellar object 3I/ATLAS, using a nucleus extraction technique on Hubble Space Telescope (HST) observations taken between December 2025 and January 2026. The product of the V-band geometric albedo, $p_{V}$, with the physical cross-section of the nucleus is $0.22\pm 0.07$ km2, which corresponds to an effective radius of $1.3\pm 0.2$ km if assuming a comet-like albedo $p_{V}=0.04$. This size is in agreement with an independent estimate based on the reported nongravitational acceleration and activity of the interstellar object. If the measured photometric variations are solely due to the rotation of an aspherical nucleus, the axis ratio must be $2:1$ or greater, and the rotation period $\gtrsim\!1$ hr. Leveraging the range of covered phase angles, we identified a significant opposition surge of $\sim\!0.2$ mag with a width of $3\arcdeg\pm 1\arcdeg$, which may include concurrent contributions from orbital plane crossing and tail projection, and determined a linear phase slope of $0.026\pm 0.006$ mag degree-1 for the coma dust. Compared to the preperihelion brightening trend, 3I faded more rapidly on the outbound leg, following an activity index of $4.5\pm 0.3$, not unusual in the context of solar system comets. This activity asymmetry is further corroborated by a postperihelion coma surface brightness profile that is significantly shallower than its preperihelion counterpart. From discovery statistics, we infer that multiple interstellar objects resembling 3I probably went undetected prior to the discovery of 1I/‘Oumuamua, unless the overall population possesses a steep size distribution.
@@ I Introduction
3I/ATLAS, formerly designated C/2025 N1 (ATLAS), was discovered using the Asteroid Terrestrial-impact Last Alert System (ATLAS) sky survey on 1 July 2025. Serendipitous prediscovery observations were quickly identified (e.g. A. D. Feinstein et al., 2025; J. Martinez-Palomera et al., 2025; Q. Ye et al., 2025), along with numerous followup observations, leading to the definitive identification of the object as the third interstellar object known to intrude into our solar system (L. Denneau et al., 2025). Its trajectory is significantly more hyperbolic with respect to the Sun than those of the previous two interstellar objects, 1I/‘Oumuamua and 2I/Borisov, with an eccentricity of $e=6.14$. The closest approach of the object to the Sun occurred on 29 October 2025 at a perihelion distance of $q=1.4$ au, implying an excess speed of $\sim\!58$ km s-1, which suggests a kinematic age of 3-11 Gyr for 3I, older than the previous interstellar objects (A. G. Taylor & D. Z. Seligman, 2025). The object is statistically unlikely to have approached any star as closely as the current approach to the Sun (D. Jewitt & J. Luu, 2025).
Numerous observations of 3I have been conducted using both ground-based facilities and spacecraft, spanning a wide range of wavelengths and utilising various techniques including imaging, spectroscopy, and polarimetry. In brief the interstellar object was reported to exhibit a steady growth of activity as it approached the Sun (e.g. M. R. Combi et al., 2025; D. Jewitt & J. Luu, 2025; T. Santana-Ros et al., 2025), with peculiar dust scattering properties (Z. Gray et al., 2025; S. Choi et al., 2026), slightly enriched volatile content, and distinct outgassing behaviour (e.g. M. A. Cordiner et al., 2025; D. Hutsemékers et al., 2025; N. X. Roth et al., 2025). Its activity was dominated by CO2 at heliocentric distances $r_{\rm H}\gtrsim 3$ au, which was later overtaken by H2O sublimation when closer to the Sun (e.g. M. R. Combi et al., 2025; M. A. Cordiner et al., 2025; Z. Xing et al., 2025; B. Yang et al., 2025). However, there has been no successful observation of the nucleus of 3I to date, primarily because of hindrance by its activity, which began prior to the earliest available observations (e.g. Q. Ye et al., 2025). Our previous attempt to extract the nucleus signal of 3I from preperihelion Hubble Space Telescope (HST) observations in July 2025 yielded an upper limit of $\sim\!2.8$ km for the nucleus radius (D. Jewitt et al., 2025). Nevertheless, characterising the radius of the nucleus remains scientifically pivotal, as an essential input for physical models and a constraint on the population of interstellar objects as a whole. In this letter, we report the first detection of the nucleus of 3I using postperihelion optical imaging from our HST program.
@@ III.1 Conservative Constraint
The apparent magnitudes are functions of heliocentric distance ($r_{\rm H}$), HST-centric distances ($\Delta$), and phase angle ($\alpha$), all of which evolved over the course of our HST observations. We normalised both distances to $r_{\oplus}=1$ au and corrected for phase effects by assuming a linear phase model with a slope of $\beta_{\alpha}=0.04\pm 0.02$ mag deg-1 (e.g. P. L. Lamy et al., 2004; M. M. Knight et al., 2024) to compute the absolute $V$-band magnitude of the nucleus:
@@
where $m_{{\rm n},V}$ is the apparent $V$-band magnitude of the nucleus, $r_{\rm H}$ and $\Delta$ are expressed in au, and $\alpha$ is phase angle in degrees. We thereby determined a lower bound of $H_{{\rm n},V}\gtrsim 15$ across the five HST epochs for the nucleus of the interstellar object, consistent with $H_{{\rm n},V}\gtrsim 15.4$ from the preperihelion HST measurement (D. Jewitt et al., 2025).
@@ III.2 Nucleus Extraction
$\Sigma\left(\bm{\rho}\right)$ $=\left[k_{\rm n}\delta\left(\bm{\rho}\right)+k_{\rm c}\left(\theta\right)\left(\dfrac{\rho}{\rho_{0}}\right)^{-\gamma\left(\theta\right)}\right]\ast\mathcal{P}\left(\bm{\rho}\right)$
$=k_{\rm n}\mathcal{P}\left(\bm{\rho}\right)+\left[k_{\rm c}\left(\theta\right)\left(\dfrac{\rho}{\rho_{0}}\right)^{-\gamma\left(\theta\right)}\right]\ast\mathcal{P}\left(\bm{\rho}\right),$ (3)
where the observed total surface brightness, $\Sigma$, is a function of the projected vector from the photocenter, $\bm{\rho}$, having length $\rho$ and azimuthal angle $\theta$. The intrinsic nucleus signal is modelled as a two-dimensional Dirac delta function $\delta$, scaled by a constant factor $k_{\rm n}$. The coma contribution is parametrised by scaling factor $k_{\rm c}$ and logarithmic surface brightness gradient $\gamma$, both as functions of the azimuthal angle. A spherically symmetric coma in steady state has $\gamma=1$, whereas solar radiation pressure steepens the gradient to $\gamma=1.5$ in the antisunward direction (D. C. Jewitt & K. J. Meech, 1987). The observed signal is the result of convolving the intrinsic profile with the WFC3 camera PSF ($\mathcal{P}$), which we simulated using the TinyTim package (J. E. Krist et al., 2011). In Equation (3) the symbol $\ast$ denotes the convolution operator. We adopted a normalisation distance of $\rho_{0}=1$ pixel to ensure that the scaling factors for the nucleus and the coma share the same units.
@@
We next performed PSF photometry on the residual signal with psffit from Buie’s IDL library. Results for each HST visit are listed in Table 2, while comparisons of the observed and modelled radial brightness profiles and the apparent magnitude time series are shown in Figures 3 and 4, respectively. We were fully aware that the extracted nucleus flux possibly fell inside the caution zone where results can be biased by the nucleus extraction technique (M.-T. Hui & J.-Y. Li, 2018). To investigate the reliability of the results, we varied the azimuthal smoothing window width, image subsampling factor, and the fitted annular region within the ranges suggested by M.-T. Hui & J.-Y. Li (2018), and we even switched to aperture photometry on the residual feature using an aperture radius corresponding to its FWHM so as to better exclude jet-like features in the near-nucleus region. We found the results to be consistent within the $1\sigma$ significance level regardless of the parameter choices or photometry method. Furthermore, applying the method described in Section III.1 to the whole image sequence yielded the same temporal trends as those shown in Figure 4, albeit with a systematic brightness offset due to contamination from the coma. Therefore, we conclude that the nucleus extraction technique successfully revealed the nucleus of 3I. Our previous attempt using preperihelion HST observations (D. Jewitt et al., 2025) was likely unsuccessful both as a result of lower spatial resolution resulting from a greater (3.0 au) geocentric distance and a steeper preperihelion radial brightness profile of the coma.
Figure 4 shows marginal evidence for brightness variability, consistent with the rotation of an aspherical nucleus and/or with residual near-nucleus coma imperfectly removed by the extraction technique. Unfortunately, our data are too sparsely sampled to permit any useful estimate of the nucleus rotation period, except that it is likely $\gtrsim\!1$ hour. Under the assumption that the brightness variations are real and driven purely by rotation, the maximum range of the apparent magnitude within a single HST epoch, $\Delta m_{{\rm n},V}\gtrsim 0.8$ mag, implies a nucleus axis ratio projected into the plane of the sky of $10^{0.4\Delta m_{{\rm n},V}}\gtrsim 2:1$. However, this axis ratio would be lower, even potentially approaching unity, if a significant portion of the variability stems from near-nucleus dust activity.
Averaging these measurements yields an absolute $V$-band magnitude of $H_{{\rm n},V}=17.1\pm 0.4$ for the nucleus of 3I, corresponding to an effective optical cross-section of $C_{\rm n}=0.22\pm 0.07$ km2. Assuming a $V$-band geometric albedo of $p_{V}=0.04$, the effective nucleus radius is $R_{\rm n}=1.3\pm 0.2$ km, much larger than the nuclei of 1I/‘Oumuamua (highly aspherical, with an effective radius of $\sim\!0.08$ km) and 2I/Borisov ($\sim\!0.4$ km) (D. Jewitt & D. Z. Seligman, 2023, and citations therein). In Figure 4 we denote the predicted apparent magnitude of a spherical nucleus having the same effective mean radius with a blue dashed line to compare it with the time sequence of individual apparent magnitude measurements. The shaded area represents the $1\sigma$ uncertainty region propagated from the standard deviation of the mean and the assumed error in the adopted linear phase function.
@@
Figure 2: Extraction of the nucleus of 3I/ATLAS. The observed HST image (first panel from the left) is shown alongside the best-fit coma model (second panel) for each HST visit (indicated at right). Subtraction of the model from the observation reveals the nucleus of 3I (third panel). For illustrative purposes, the displayed images are median-combined from individual exposures within each visit and displayed with a logarithmic intensity stretch. In each row, the red and magenta arrows indicate local north and east, respectively, in the J2000 equatorial coordinate system, with the projected antisolar direction and the negative heliocentric velocity of 3I represented by the yellow and cyan arrows, respectively. The horizontal white bar near the bottom marks a scale of 1″ in apparent length.
@@
Figure 4: Temporal apparent magnitude variations of the nucleus of 3I across the HST visits. In each panel, the horizontal red dashed line marks the mean value computed from the specific visit. The blue dashed line represents the predicted apparent magnitude of a spherical nucleus of the same effective radius, with its corresponding $1\sigma$ uncertainty region shaded in grey.
Table 3: Comparisons of Nucleus Size Estimates of 3I/ATLAS footnotetext: The collimation coefficient of activity was assumed to be $\kappa=0.5$. For alternative values of $\kappa\in\left(0,1\right)$, the size estimate scales as $\propto\kappa^{1/3}$ per Equation (4). Note that the reported formal error does not take into account the systematic uncertainty in $\kappa$, and hence it is only an underestimate of the actual uncertainty.footnotetext: To ensure consistency, estimates originally derived by other authors assuming $p_{V}=0.05$ have been rescaled to match our adopted value of $p_{V}=0.04$.
@@ IV.1 Nucleus Size Inferred from Nongravitational Effect
The nongravitational acceleration of 3I enables an independent estimate of its nucleus size, which can then be compared with the results derived from our HST observations. The orbital solution of 3I by JPL Horizons demonstrates a statistically significant detection of nongravitational accelerations, presumably due to anisotropic mass loss. Adopting an inverse-square law for the nongravitational force model, i.e. $g_{\rm NG}\propto r_{\rm H}^{-2}$, the solution yields the radial, transverse, and normal (RTN) components of the nongravitational parameters, originally defined by B. G. Marsden et al. (1973), to be $A_{1}=\left(+4.5\pm 0.1\right)\times 10^{-8}$ au day-2, $A_{2}=\left(+1.7\pm 0.1\right)\times 10^{-8}$ au day-2, and $A_{3}=\left(-6.0\pm 0.2\right)\times 10^{-9}$ au day-2, respectively. These components all exhibit a significance of $\gtrsim\!10\sigma$. [^2] Retrieved on 5 February 2026. Based on the conservation of momentum and assuming a spherical nucleus, the nucleus radius can be inferred from the nongravitational effect and the outgassing mass-loss rate using the formulation by M.-T. Hui et al. (2020):
@@
We evaluated the nongravitational effect of 3I around the time of perihelion when its outgassing activity and the resulting nongravitational acceleration were most pronounced as a consequence of momentum conservation. The benefit of picking an epoch around perihelion is that the dominance of water production in the outgassing activity of 3I over other sublimating species, including CO2, allowed us to safely neglect contributions from these various substances, which can no longer be omitted at greater heliocentric distances. At $r_{\rm H}=1.401$ au, the interstellar object exhibited significant water production at a rate of $Q_{\rm gas}=\left(3.2\pm 1.0\right)\times 10^{29}$ s-1, which subsequently declined during the outbound leg (M. R. Combi et al., 2025).  [^3] At the time of writing, H. Tan et al. (2026) reported water production rates up to an order of magnitude lower than M. R. Combi et al. (2025), despite using the same dataset. While this difference deserves to be investigated, applying the correlation from L. Jorda et al. (2008) to the near-perihelion magnitudes of 3I reported by Q. Zhang & K. Battams (2025) yielded water production rates of 3I in good agreement with M. R. Combi et al. (2025) at similar heliocentric distances, albeit preperihelion, and so we do not adopt the results of H. Tan et al. (2026). Setting $\mathscr{U}_{\rm gas}=18$ for water, $\kappa=0.5$ to approximate hemispherical emission, and adopting the empirical law for outflow speed of gas from A. L. Cochran & D. G. Schleicher (1993) yielding $v_{\rm gas}\approx 0.7$ km s-1, we obtained the nucleus radius of 3I from Equation (4) to be $1.5\pm 0.1$ km, in excellent agreement with our results in Section III.2 via the nucleus extraction technique, thereby solidifying our nucleus size estimate.
@@
Here the formal uncertainty was properly propagated from the covariance matrix of the RTN nongravitational parameters provided by Horizons and the reported error in the water production rate in M. R. Combi et al. (2025). However, this formal uncertainty is a lower limit to the true uncertainty because it neglects systematic effects. The maximum nucleus size permitted by the detected nongravitational acceleration occurs under a fully collimated mass-loss scenario, i.e. $\kappa=1$, which yields $R_{\rm n}\leq 1.9\pm 0.2$ km. Conversely, a much smaller nucleus size estimate will be inferred if $\kappa$ is much smaller, corresponding to less anisotropic mass loss. Furthermore, if a significant fraction of the measured water production is due to the sublimation of icy grains in the coma rather than from the nucleus surface, our size estimates inferred from the nongravitational effect should be considered as upper limits to the actual nucleus radius. We cannot reliably quantify the systematic uncertainties given the presently available data.
@@
In Table 3, we compare our nucleus size estimates of 3I with earlier results. Our inferred nucleus radius is large compared to earlier results by T. M. Eubanks et al. (2025) and J. C. Forbes & H. Butler (2025), but small compared to that of R. Cloete et al. (2025), despite all three groups applying the same methodology. We posit that earlier solutions for the nongravitational parameters were less reliable due to the shorter observed arc, where potential biases, such as tailward offsets in astrometric measurements, could be more easily absorbed into the orbital fits. Furthermore, the applied nongravitational force model might have been overly simplified and therefore failed to closely approximate earlier epochs when the outgassing of 3I was still weak and the nongravitational effect was far less significant. For instance, at greater heliocentric distances, the nongravitational acceleration could have been arisen from outgassing of multiple species rather than solely from the dominant volatile. The discrepancy with R. Cloete et al. (2025) stems from their reliance on an early upper limit for the nongravitational acceleration, which is inconsistent with the more recent Horizons solution by nearly two orders of magnitude, and to a lesser extent from their neglect of the collimation factor, $\kappa$.